OpenStudy (anonymous):

Explain how the formula for the area of a regular polygon is derived from that of a triangle.

OpenStudy (anonymous):

take for example a square which is a regular polygon if we divide the square in two congruent triangles then A of square = 1/2( b x h ) + 1/2 (b x h) A = 1/2 (s x s) + 1/2 (s x s) = 1/2 S^2 + 1/2 S^2 = S^2